# RRADistill: Distilling LLMs’ Passage Ranking Ability for Document Re-Ranking of Long-Tail Queries in a Search Engine

### v1

@@ Abstract
for Document Re-Ranking of Long-Tail Queries in a Search Engine
Nayoung Choi2$\dagger$*, Youngjune Lee1$\dagger$, Gyu-Hwung Cho1, Haeyu Jeong1, Jungmin Kong1, Saehun Kim1, Keunchan Park1, Jaeho Choi1, Sarah Cho1, Inchang Jeong1, Gyohee Nam1, Sunghoon Han1, Wonil Yang1 1Naver Corporation, 2Emory University nayoung.choi@emory.edu, youngjune.lee93@navercorp.com
@@ 1 Introduction
Large Language Models (LLMs), such as ChatGPT OpenAI (2022) and GPT-4 Achiam et al. (2023), have shown remarkable potential across diverse search tasks, including query rewriting Mao et al. (2023); Dhole and Agichtein (2024), query and document expansions Wang et al. (2023a); Mackie et al. (2023); Ma et al. (2023a). As LLMs advance in complex tasks, they also show potential for passage ranking, which involves understanding relationships between queries and multiple documents. Recent studies Qin et al. (2024); Ma et al. (2023b); Zhuang et al. (2024); Pradeep et al. (2023a) show that instruction-tuned LLMs like GPT-4, Flan-T5 Chung et al. (2024) and Vicuna Chiang et al. (2023) effectively handle passage ranking in zero-shot settings. Motivated by these studies, we also conducted zero-shot re-ranking with our in-house LLMs, HyperCLOVA X (HCX) Yoo et al. (2024) which comprises Large (HCX-L) and Small (HCX-S) variants. We put a query and a list of document snippet texts to HCX, using a Korean translation of list-wise prompt from RankGPT Sun et al. (2023). We found that HCX-L effectively returns document identifiers in the desired order, excluding low-relevance ones, as depicted in Figure 1. Unlike short-head queries, which are popular and short like keywords, long-tail queries are complex and involve longer, specific phrases (A.1.1). These queries benefit more from relevance than feedback, and semantic than syntactic matching due to their rich semantic content but lack of user feedback. Thus, LLMs’ ability to rank complex long-tail queries by relevance is highly valuable.
@@
However, the slow inference speed challenges the direct use of LLMs in search engine. To address this, we trained a much smaller Language Model (SLM) to retain HCX-L’s ranking ability, following a trend known as LLM distillation, similar to RankGPT and TWOLAR Baldelli et al. (2024). This involves two stages: 1) Generating ranking label using LLMs, and 2) Training the SLM ranker. When generating ranking label with LLMs, previous studies utilize the list-wise permutation generation method, which inputs a query and a set of documents to LLM and receives an ordered list of document identifiers. However, these approach require sliding windows, which infer multiple times on partial lists due to the prompt length constraints of LLMs, causing a burden. Moreover, previous studies viewed the missing phenomenon as a problem, where LLMs fail to include all input documents in the output. However, we observed that in most cases, excluded documents due to missing are significantly irrelevant to the query, making it a valuable signal. Consequently, we reframed the missing and highlighted its impact. We developed our own label generation pipeline to address these issues, including two key techniques: 1) Pre-rank to filter documents, retaining only those effective to train SLM rankers and bypass the sliding window, 2) Consider missing as useful signals, and utilize excluded documents as hard negatives, to train SLM rankers. Our pipeline speeds up labeling and provides compact yet effective training data.
@@ 4 Conclusion
In this paper, we present an efficient label generation pipeline using LLMs that utilizes a pre-ranker to select effective documents and leverages LLMs’ missing phenomenon as a useful signal. We also present effective training methods to capture long and complex query-document relevance in both BERT and GPT. However, only key parts of the model structure are employed during inference, minimizing the service deployment burden. Through extensive experiments, including A/B testing on NAVER Search, we demonstrate the effectiveness of our approach for long-tail queries re-ranking.

### v3

# RRADistill: Distilling LLMs’ Passage Ranking Ability for Long-Tail Queries Document Re-Ranking on a Search Engine

@@ Abstract
for Long-Tail Queries Document Re-Ranking on a Search Engine
Nayoung Choi2$\dagger$*, Youngjune Lee1$\dagger$, Gyu-Hwung Cho1, Haeyu Jeong1, Jungmin Kong1, Saehun Kim1, Keunchan Park1, Sarah Cho1, Inchang Jeong1, Gyohee Nam1, Sunghoon Han1, Wonil Yang1 and Jaeho Choi1 1Naver Corporation, 2Emory University nayoung.choi@emory.edu, youngjune.lee93@navercorp.com
@@ 1 Introduction
Providing relevant search results to enhance user satisfaction is crucial in search engines. To achieve this, various research efforts have been conducted on different search tasks, such as retrieval and ranking Formal et al. (2021); Khattab and Zaharia (2020); Nogueira and Cho (2019); Kang et al. (2024a, b). Large Language Models (LLMs), such as ChatGPT OpenAI (2022) and GPT-4 Achiam et al. (2023), have shown remarkable potential across diverse search tasks, including query rewriting Mao et al. (2023); Dhole and Agichtein (2024), query and document expansions Wang et al. (2023a); Mackie et al. (2023); Ma et al. (2023a). As LLMs advance in complex tasks, they also show potential for passage ranking, which involves understanding relationships between queries and multiple documents. Recent studies Qin et al. (2024); Ma et al. (2023b); Zhuang et al. (2024); Pradeep et al. (2023a) show that instruction-tuned LLMs like GPT-4, Flan-T5 Chung et al. (2024) and Vicuna Chiang et al. (2023) effectively handle passage ranking in zero-shot settings. Motivated by these studies, we also conducted zero-shot re-ranking with our in-house LLMs, HyperCLOVA X (HCX) Yoo et al. (2024) which comprises Large (HCX-L) and Small (HCX-S) variants. We put a query and a list of document snippet texts to HCX, using a Korean translation of list-wise prompt from RankGPT Sun et al. (2023). We found that HCX-L effectively returns document identifiers in the desired order, excluding low-relevance ones, as depicted in Figure 1. Unlike short-head queries, which are popular and short like keywords, long-tail queries are complex and involve longer, specific phrases (A.1.1). These queries benefit more from relevance than feedback, and semantic than syntactic matching due to their rich semantic content but lack of user feedback. Thus, LLMs’ ability to rank complex long-tail queries by relevance is highly valuable.
@@
However, the slow inference speed challenges the direct use of LLMs in search engine. To address this, we trained a much smaller Language Model (SLM) to retain HCX-L’s ranking ability. This area of research is known as knowledge distillation Lee and Kim (2021); Kang et al. (2020, 2021); Lee et al. (2024); Kang et al. (2023), where the knowledge of large models is transferred to smaller models to improve efficiency in information retrieval. Our approach builds on similar LLM distillation methods, such as those used in RankGPT and TWOLAR Baldelli et al. (2024). This involves two stages: 1) Generating ranking label using LLMs, and 2) Training the SLM ranker. When generating ranking label with LLMs, previous studies utilize the list-wise permutation generation method, which inputs a query and a set of documents to LLM and receives an ordered list of document identifiers. However, these approach require sliding windows, which infer multiple times on partial lists due to the prompt length constraints of LLMs, causing a burden. Moreover, previous studies viewed the missing phenomenon as a problem, where LLMs fail to include all input documents in the output. However, we observed that in most cases, excluded documents due to missing are significantly irrelevant to the query, making it a valuable signal. Consequently, we reframed the missing and highlighted its impact. We developed our own label generation pipeline to address these issues, including two key techniques: 1) Pre-rank to filter documents, retaining only those effective to train SLM rankers and bypass the sliding window, 2) Consider missing as useful signals, and utilize excluded documents as hard negatives, to train SLM rankers. Our pipeline speeds up labeling and provides compact yet effective training data.
@@ 4 Conclusion
In this paper, we present an efficient label generation pipeline using LLMs that utilizes a pre-ranker to select effective documents and leverages LLMs’ missing phenomenon as a useful signal. We also present effective training methods to capture long and complex query-document relevance in both BERT and GPT. However, only key parts of the model structure are employed during inference, minimizing the service deployment burden. Through extensive experiments, including A/B testing on NAVER Search, we demonstrate the effectiveness of our approach for long-tail queries re-ranking. We believe that our approach, which captures context while focusing on key terms or features, will perform effectively across various domains, such as e-commerce Lee et al. (2023); Wang et al. (2022); Kang et al. (2022). We expect that it will serve as a valuable direction for future applications in industry.